# Scalable learning of macroscopic stochastic dynamics

### v1

@@ Abstract
Macroscopic dynamical descriptions of complex physical systems are crucial for understanding and controlling material behavior. With the growing availability of data and compute, machine learning has become a promising alternative to first-principles methods to build accurate macroscopic models from microscopic trajectory simulations. However, for spatially extended systems, direct simulations of sufficiently large microscopic systems that inform macroscopic behavior is prohibitive. In this work, we propose a framework that learns the macroscopic dynamics of large stochastic microscopic systems using only small-system simulations. Our framework employs a partial evolution scheme to generate training data pairs by evolving large-system snapshots within local patches. We subsequently identify the closure variables associated with the macroscopic observables and learn the macroscopic dynamics using a custom loss. Furthermore, we introduce a hierarchical upsampling scheme that enables efficient generation of large-system snapshots from small-system trajectory distributions. We empirically demonstrate the accuracy and robustness of our framework through a variety of stochastic spatially extended systems, including those described by stochastic partial differential equations, idealised lattice spin systems, and a more realistic NbMoTa alloy system.
1 Introduction
@@
To obtain accurate time evolution of macroscopic observables, large-scale microscopic simulations over extended times are often required, yet their high computational expense remains the main bottleneck [2, 3, 4]. For example, Density Functional Theory (DFT) and the related Ab Initio Molecular Dynamics (AIMD) represent a milestone in computational methods for studying molecules and solid-state materials at the quantum mechanical level. However, due to the well-known exponential wall challenge, where computational expense increases rapidly with the number of particles, DFT simulations are usually limited to relatively small systems that may be insufficient in capturing true macroscopic behavior [5, 6].
@@
Figure 1: Schematic illustration of our framework. The hierarchical upsampling scheme generates the large-system distribution $\mathcal{D}$ from the small-system trajectory distribution $\mathcal{D}_{s}$ through multiple iterations, each consisting of an Upsample and a LocalRelax step. An example of one iteration for the Ising model is shown. For the partial evolution scheme, for every $\mathbf{x}_{t}\sim\mathcal{D}$, a patch $\mathcal{I}$ is first uniformly sampled, then the microscopic dynamics is evolved locally within the patch $\mathcal{I}$ for a short time to yield $\mathbf{x}_{t+\delta t,\mathcal{I}}$. For the closure modeling, an autoencoder is trained to discover the closure variables to the macroscopic observables, and the macroscopic dynamics are identified with the designed loss $\mathcal{L}_{p}$.
@@
Various approaches have been developed to overcome the computational bottleneck. The Kinetic Monte Carlo (KMC) algorithm represents the microscopic dynamics as a Markov chain by coarse-graining the time axis [7, 8, 9]. The transition rates for all possible events need to be computed for every KMC step. Therefore, KMC is still computationally prohibitive for large systems [10, 11]. Machine learning force fields (MLFFs) replace the expensive ab initio force computations with efficient neural network–based predictions [12, 13, 14, 15]. When applying MLFFs to molecular dynamics simulations, the time step of the molecular dynamics simulation must be chosen on the scale of femtoseconds to capture atomic vibrations. Consequently, millions to billions of integration steps are required for the simulation, and atomic forces are computed for each time step, still resulting in high computational cost for large systems [16]. Coarse-grained MLFFs further improve computational efficiency by mapping several atoms onto effective particles, thereby reducing the number of degrees of freedom. However, the training of coarse-grained MLFFs requires microscopic simulation data, and the largest barrier to applying coarse-grained MLFFs to large systems is generating enough training data [17, 18, 19, 20, 21]. The closure modeling methods, instead, model the dynamics of macroscopic observables directly, but still rely on short microscopic simulations of the large system or microscopic forces on all atoms for macroscopic dynamics identification of large systems [22].
@@
Despite their methodological differences, existing approaches all rely either on direct microscopic simulations or on training data derived from such simulations. However, due to the computational constraint, microscopic simulation for large systems with millions to billions of atoms over extended time is generally intractable. This leads to a natural question: Can accurate macroscopic dynamics of large systems be obtained when only small-system microscopic simulations are accessible? Chen and Li [23] proposed a training procedure on the microscopic coordinates that addresses this question for deterministic dynamics. However, their method requires partial computation of microscopic forces for the macroscopic dynamics identification. In the case of stochastic systems, where the dynamics are described by the conditional distribution of the next configuration given the current one, such microscopic forces are generally not well-defined. Hence, their method cannot be easily generalized to stochastic systems. Yet, stochastic microscopic systems are arguably more prevalent, especially in the modeling of chemical reactions, molecular dynamics, and ferromagnetic phase transitions [24]. The main goal of our method is to address the above question for stochastic dynamical systems.
@@
In this work, we develop a framework that can accurately identify the macroscopic dynamics of stochastic microscopic systems, while requiring only microscopic simulations of small systems. Specifically, given the data distribution $\mathcal{D}$, composed of snapshots of large systems spanning states from far-from-equilibrium to near-equilibrium, we introduce a partial evolution scheme which evolves $\mathbf{x}_{t}\sim\mathcal{D}$ locally within a local patch $\mathcal{I}$ for a short time $\delta t$ to produce training data pairs $\{\mathbf{x}_{t},\mathbf{x}_{t+\delta t,\mathcal{I}}\}$. Next, building on the workflow of Ref. [22], we identify the closure variables to the macroscopic observables and model the resulting dynamics via stochastic differential equations (SDE). To adapt to the training data generated through the partial evolution scheme, we introduce a modified SDE loss and provide a theoretical justification. In addition, we design a hierarchical upsampling scheme that efficiently generates $\mathcal{D}$ from the small-system trajectory distribution $\mathcal{D}_{s}$.
@@
The key idea of our framework is summarized in Fig.˜1. We provide a detailed description of each component in Section˜2. We further validate the accuracy and robustness of our method through a variety of stochastic microscopic systems, including stochastic partial differential equations (SPDE) systems, spin systems, and a more realistic NbMoTa alloy system in Section˜3.
2 METHODOLOGY
@@
In many real applications, we are interested in the dynamics of some macroscopic observables, denoted by $\mathbf{z}^{\ast}=\boldsymbol{\varphi}^{\ast}(\mathbf{x})$. The form of $\boldsymbol{\varphi}^{\ast}$ is given beforehand, and $\boldsymbol{\varphi}^{\ast}$ can be applied to different system sizes. Since the underlying microscopic system is spatially extended, we are interested in the intensity quantities that do not scale with system size. For instance, in the Ising model, it is common to study the average magnetization $M=\sum_{i=1}^{n}\mathbf{x}_{i}/n$ instead of the total magnetization. In what follows, we will limit our discussion of macroscopic observables to intensity quantities.
Assume we are given a microscopic simulator $\mathcal{S}_{n_{s}}$, which can accurately simulate the microscopic dynamics of a small system up to size $n_{s}\ll n$ due to computational constraints. From this simulator, we obtain the trajectory distribution $\mathcal{D}_{s}$ of the small system. The goal of this work is to identify the macroscopic dynamics of a large system of size $n$ using only such small-scale simulations.
2.1 Closure modeling of macroscopic dynamics
Existing works on macroscopic dynamics identification typically involve two components: discovering the closures for macroscopic observables, and jointly identifying their dynamics [22, 25, 26]. Our work adopts the same two components for macroscopic dynamics identification. Assume we are given the data distribution $\mathcal{D}$ of the large system, consisting of multiple snapshots of the large system, we will first generate temporal training data pairs as follows.
@@ 2.1 Closure modeling of macroscopic dynamics
We partition the underlying regular lattice into $K=n/n_{s}$ small patches, each containing $n_{s}$ lattice sites. For example, for the two-dimensional Ising model on a $64^{2}$ square lattice, we partition the large square into $64$ square patches of size $n_{s}=8^{2}$. Let the index set of lattice sites in the $k$-th patch be $\mathcal{I}^{k}$, which is a subset of $\{1,\cdots n\}$ and contains $n_{s}$ lattice sites. The state of the lattice sites in patch $\mathcal{I}^{k}$ is then written as $\mathbf{x}_{\mathcal{I}^{k}}=\{\mathbf{x}_{i}\}_{i\in\mathcal{I}^{k}}$, and the microscopic state as $\mathbf{x}=(\mathbf{x}_{\mathcal{I}^{1}},\cdots,\mathbf{x}_{\mathcal{I}^{K}})$.
For each configuration $\mathbf{x}_{t}\sim\mathcal{D}$, we uniformly sample a patch $\mathcal{I}$ from the $K$ patches with probability $1/K$. Next, the microscopic simulator $\mathcal{S}_{n_{s}}$ is used to evolve $\mathbf{x}_{t}$ within the selected patch $\mathcal{I}$ for a short time $\delta t$, yielding the updated state $\mathbf{x}_{t+\delta t,\mathcal{I}}$. Consequently, the resulting training data pair is $\{\mathbf{x}_{t},\mathbf{x}_{t+\delta t,\mathcal{I}}\}$. Note that the time step $\delta t$ may also be random. For instance, $\delta t$ is sampled from an exponential distribution in the case of kinetic Monte Carlo dynamics.
We will demonstrate how to identify the macroscopic dynamics from the training data pairs obtained from the partial evolution scheme.
@@
Once the autoencoder is trained, we will generate the latent training data pair $\{\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}}\}$ for the macroscopic dynamics identification:
@@
where $\mathbf{x}_{t,\mathcal{I}}$ denotes the restriction of $\mathbf{x}_{t}$ to the local patch $\mathcal{I}$. Next, we introduce the process of identifying macroscopic dynamics.
Macroscopic dynamics identification
@@
where $\boldsymbol{\mu}$ is the drift term and $\boldsymbol{\Sigma}$ is the diffusion term. In most experiments, we adopt fully connected networks for both $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$. For the NbMoTa alloy experiment in Section˜3.4, we adopt OnsagerNet [27, 22] for $\boldsymbol{\mu}$ to account for more complex macroscopic dynamics. The OnsagerNet can capture physically interpretable and stable macroscopic dynamics by incorporating the generalized Onsager principle into the model structure, and its form is given by:
$\boldsymbol{\mu}(\mathbf{z}_{t})=-(\mathbf{M}(\mathbf{z}_{t})+\mathbf{W}(\mathbf{z}_{t})))\nabla V(\mathbf{z}_{t})+\mathbf{f}(\mathbf{z}_{t}),$ (5)
@@
Existing works train the SDE by minimize the negative log-likelihood [22, 28, 29, 30]:
$\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]=\mathbb{E}_{\mathbf{z}_{t},\mathbf{z}_{t+\delta t}}[-2\log p\left(\mathbf{z}_{t+\delta t}\,\big|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t\right)],$ (6)
@@
$\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]=\mathbb{E}_{\mathbf{z}_{t},\mathbf{z}_{t+\delta t,{\color[rgb]{1,0,0}\mathcal{I}}}}[-2\log p(\mathbf{z}_{t+\delta t,{\color[rgb]{1,0,0}\mathcal{I}}}\,|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,{\color[rgb]{1,0,0}K}\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t)].$ (7)
The only difference between $\mathcal{L}$ and $\mathcal{L}_{p}$ is that we multiply a coefficient $K$ to the diffusion term in $\mathcal{L}_{p}$. We can interpret the influence of $K$ qualitatively. Since $K>1$ scales the diffusion term, $K$ will reduce the magnitude of the diffusion term. During the data generation of $\mathbf{x}_{t+\delta t,\mathcal{I}}$ from $\mathbf{x}_{t}$, we introduce additional randomness by performing partial evolution. Therefore, in the identification of macroscopic dynamics, we multiply the diffusion term by $K$ to correct for the extra stochasticity.
@@
It is important to observe that in Eq.˜3, we define $\mathbf{z}_{t+\delta t,\mathcal{I}}$ in a nontrivial way, while the most naive way will be $\mathbf{z}_{t+\delta t,\mathcal{I}}=\varphi(\mathbf{x}_{t+\delta t,\mathcal{I}})$. Note that the loss function $\mathcal{L}_{p}(\boldsymbol{\theta})$ involves the weighted norm of the residual $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}-\boldsymbol{\mu}(\mathbf{z}_{t})\delta t$ since $p$ represents a gaussian distribution. If we define $\mathbf{z}_{t+\delta t,\mathcal{I}}=\varphi(\mathbf{x}_{t+\delta t,\mathcal{I}})$ in the naive way, we will have:
@@
which is the average of $\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})-\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{J}})$. Since $\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})$ may be very different from $\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{J}})$ when $\mathcal{J}\neq\mathcal{I}$, the resulting $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}$ will be very noisy. However, the definition in Eq.˜3 will yield:
@@
which directly measures the change in the patch $\mathcal{I}$, thus reducing noise and leading to more stable training. We refer to the approach that adopts the naive formulation $\mathbf{z}_{t+\delta t,\mathcal{I}}=\varphi(\mathbf{x}_{t+\delta t,\mathcal{I}})$ and identifies the macroscopic dynamics via $\mathcal{L}$ as the baseline. A detailed comparison of our method with the baseline is provided in Section˜3.
@@
In many microscopic systems, the microscopic interactions between lattice sites are local. For instance, microscopic interactions are limited to the first nearest neighbor in the Ising model, and are limited to a finite cutoff distance in alloy systems. Under such cases, when $\delta t$ is small enough, the increments:
$\mathbf{z}_{t+\delta t,\mathcal{J}}-\mathbf{z}_{t}$ $=\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{J}})-\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{J}}),$
can be approximately independent for disjoint patches $\mathcal{I}\neq\mathcal{J}$.
Let $\tilde{\mathbf{x}}_{t+\delta t,\mathcal{I}}$ be the microscopic state obtained by first evolving the full system from $\mathbf{x}_{t}$ for time step $\delta t$, and subsequently restricted to the local patch $\mathcal{I}$. Under the local interaction assumption and for sufficiently small $\delta t$, the distribution of $\tilde{\mathbf{x}}_{t+\delta t,\mathcal{I}}$ can be well approximated by $\mathbf{x}_{t+\delta t,\mathcal{I}}$ obtained by partial evolution:
$\tilde{\mathbf{x}}_{t+\delta t,\mathcal{I}}\overset{d}{\approx}\mathbf{x}_{t+\delta t,\mathcal{I}},\quad\forall I\in\mathcal{I}.$ (12)
Furthermore, since we consider a sufficiently large microscopic system and the macroscopic observables are intensive quantities, the macroscopic observable of the full system can be approximated by the average of the macroscopic observable on small patches:

### v2

@@ Abstract
Macroscopic dynamical descriptions of complex physical systems are crucial for understanding and controlling material behavior. With the growing availability of data and compute, machine learning has become a promising alternative to first-principles methods to build accurate macroscopic models from microscopic trajectory simulations. However, for spatially extended systems, direct simulations of sufficiently large microscopic systems that inform macroscopic behavior are prohibitive. In this work, we propose a framework that learns the macroscopic dynamics of large stochastic microscopic systems using only small-system simulations. Our framework employs a partial evolution scheme to generate training data pairs by evolving large-system snapshots within local patches. We subsequently derive the closure variables associated with the macroscopic observables and learn the macroscopic dynamics using a custom loss. Furthermore, we introduce a hierarchical upsampling scheme that enables efficient generation of large-system snapshots from small-system snapshots. We empirically demonstrate the accuracy and robustness of our framework through a variety of stochastic spatially extended systems, including those described by stochastic partial differential equations, idealised lattice spin systems, and a more realistic NbMoTa alloy system.
preprint: APS/123-QED
I Introduction
@@
To obtain accurate time evolution of macroscopic observables, large-scale microscopic simulations over extended times are often required, yet their high computational expense remains the main bottleneck [20, 32, 51]. For example, Density Functional Theory (DFT) and the related Ab Initio Molecular Dynamics (AIMD) represent a milestone in computational methods for studying molecules and solid-state materials at the quantum mechanical level. However, due to the well-known exponential wall challenge, where computational expense increases rapidly with the number of particles, DFT simulations are usually limited to relatively small systems that may be insufficient for capturing true macroscopic behavior [55, 15].
@@
Figure 1: Schematic illustration of our framework. The hierarchical upsampling scheme generates the large-system dataset $D$ from the small-system dataset $D_{s}$ through multiple iterations, each consisting of an Upsample and a LocalRelax step. An example of one iteration for the Ising model is shown. For the partial evolution scheme, for every $\mathbf{x}_{t}\in D$, a patch $\mathcal{I}$ is first uniformly sampled, then the microscopic dynamics is evolved locally within the patch $\mathcal{I}$ for a short time to yield $\mathbf{x}_{t+\delta t,\mathcal{I}}$. For the closure modeling, an autoencoder is trained to discover the closure variables to the macroscopic observables, and the macroscopic dynamics are derived with the designed loss $\mathcal{L}_{p}$.
@@
Various approaches have been developed to overcome the computational bottleneck. The Kinetic Monte Carlo (KMC) algorithm represents the microscopic dynamics as a Markov chain by coarse-graining the time axis [50, 3, 44]. The transition rates for all possible events need to be computed for every KMC step. Therefore, KMC is still computationally prohibitive for large systems [40, 49]. Machine learning force fields (MLFFs) replace the expensive ab initio force computations with efficient neural network–based predictions [58, 4, 5, 29]. When applying MLFFs to molecular dynamics simulations, the time step of the molecular dynamics simulation must be chosen on the scale of femtoseconds to capture atomic vibrations. Consequently, millions to billions of integration steps are required for the simulation, and atomic forces are computed for each time step, still resulting in high computational cost for large systems [34]. Coarse-grained MLFFs further improve computational efficiency by mapping several atoms onto effective particles, thereby reducing the number of degrees of freedom. However, the training of coarse-grained MLFFs requires microscopic simulation data, and the largest barrier to applying coarse-grained MLFFs to large systems is generating enough training data [21, 33, 37, 54, 59]. The closure modeling methods, instead, model the dynamics of macroscopic observables directly, but still rely on short microscopic simulations of the large system or microscopic forces on all atoms for macroscopic dynamics derivation of large systems [12].
@@
Despite their methodological differences, existing approaches all rely either on direct microscopic simulations or on training data derived from such simulations. However, due to the computational constraint, microscopic simulation of large systems with millions to billions of atoms over extended time is generally intractable. This leads to a natural question: Can accurate macroscopic dynamics of large systems be obtained when only small-system microscopic simulations are accessible? Chen and Li [11] proposed a training procedure on the microscopic coordinates that addresses this question for deterministic dynamics. However, their method requires partial computation of microscopic forces for the macroscopic dynamics derivation. In the case of stochastic systems, where the dynamics are described by the conditional distribution of the next configuration given the current one, such microscopic forces are generally not well-defined. Hence, their method cannot be easily generalized to stochastic systems. Yet, stochastic microscopic systems are arguably more prevalent, especially in the modeling of chemical reactions, molecular dynamics, and ferromagnetic phase transitions [53]. The main goal of our method is to address the above question for stochastic dynamical systems.
@@
In this work, we develop a framework that can accurately derive the macroscopic dynamics of stochastic microscopic systems, while requiring only microscopic simulations of small systems. Specifically, given the dataset $D$, composed of snapshots of large systems spanning states from far-from-equilibrium to near-equilibrium, we introduce a partial evolution scheme which evolves $\mathbf{x}_{t}\in D$ locally within a local patch $\mathcal{I}$ for a short time $\delta t$ to produce training data pairs $\{\mathbf{x}_{t},\mathbf{x}_{t+\delta t,\mathcal{I}}\}$. Next, building on the workflow of Ref. [12], we derive the closure variables to the macroscopic observables and model the resulting dynamics via stochastic differential equations (SDE). To account for the additional stochasticity introduced by the random selection of patches in the partial evolution scheme, we introduce a modified SDE loss and provide a theoretical justification. In addition, we design a hierarchical upsampling scheme that efficiently generates a large-system dataset $D$ from a small-system dataset $D_{s}$, which consists of snapshots sampled from trajectories of the small system.
@@
The key idea of our framework is illustrated in Fig. 1. Key notations are summarized in Appendix A. We provide a detailed description of each component in Section II. In Section III, we further validate the accuracy and robustness of our method through a variety of stochastic microscopic systems, including stochastic partial differential equations (SPDE) systems, spin systems, and a more realistic NbMoTa alloy system.
II METHODOLOGY
@@
In many real applications, we are interested in the dynamics of some macroscopic observables, denoted by $\mathbf{z}^{\ast}=\boldsymbol{\varphi}^{\ast}(\mathbf{x})$. The form of $\boldsymbol{\varphi}^{\ast}$ is given beforehand, and $\boldsymbol{\varphi}^{\ast}$ can be applied to different system sizes. Since the underlying microscopic system is spatially extended, we are interested in the intensive quantities that do not scale with system size. For instance, in the Ising model, it is common to study the average magnetization $M=\sum_{i=1}^{n}\mathbf{x}_{i}/n$ instead of the total magnetization. In what follows, we will limit our discussion of macroscopic observables to intensive quantities.
Assume we are given a microscopic simulator $\mathcal{S}_{n_{s}}$, which can accurately simulate the microscopic dynamics of a small system up to size $n_{s}\ll n$ due to computational constraints. From this simulator, we obtain the dataset $D_{s}$ of the small system composed of snapshots sampled from trajectories of the small system. The goal of this work is to derive the macroscopic dynamics of a large system of size $n$ using only such small-scale simulations.
II.1 Closure modeling of macroscopic dynamics
Existing works on macroscopic dynamics derivation typically involve two components: discovering the closures for macroscopic observables, and jointly deriving their dynamics [12, 25, 6]. Our work adopts the same two components for macroscopic dynamics derivation. Assume we are given the dataset $D$ of the large system, consisting of multiple snapshots of the large system, we will first generate temporal training data pairs as follows.
@@ II.1 Closure modeling of macroscopic dynamics
We propose a scheme for locally evolving the microscopic dynamics of a large system within a small spatial patch, which we refer to as the partial evolution scheme. The purpose of this scheme is to generate locally evolved training data pairs $\{\mathbf{x}_{t},\mathbf{x}_{t+\delta t,\mathcal{I}}\}$ by evolving the microscopic dynamics on a small patch for a short time interval.
More specifically, we partition the underlying regular lattice into $K=n/n_{s}$ small patches, each containing $n_{s}$ lattice sites. For example, for the two-dimensional Ising model on a $64^{2}$ square lattice, we partition the large square into $64$ square patches of size $n_{s}=8^{2}$. Let the index set of lattice sites in the $k$-th patch be $\mathcal{I}^{k}$, which is a subset of $\{1,\cdots,n\}$ and contains $n_{s}$ lattice sites. The state of the lattice sites in patch $\mathcal{I}^{k}$ is then written as $\mathbf{x}_{\mathcal{I}^{k}}=\{\mathbf{x}_{i}\}_{i\in\mathcal{I}^{k}}$, and the microscopic state as $\mathbf{x}=(\mathbf{x}_{\mathcal{I}^{1}},\cdots,\mathbf{x}_{\mathcal{I}^{K}})$. For each configuration $\mathbf{x}_{t}\sim D$, we uniformly sample a patch $\mathcal{I}$ from the $K$ patches with probability $1/K$. Next, the microscopic simulator $\mathcal{S}_{n_{s}}$ is used to evolve $\mathbf{x}_{t}$ within the selected patch $\mathcal{I}$ for a short time $\delta t$, yielding the updated state $\mathbf{x}_{t+\delta t,\mathcal{I}}$. Consequently, the resulting training data pair is $\{\mathbf{x}_{t},\mathbf{x}_{t+\delta t,\mathcal{I}}\}$. We denote the resulting conditional distribution of $\mathbf{x}_{t+\delta t,\mathcal{I}}$ as $q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})$. Note that the time step $\delta t$ may also be random. For instance, $\delta t$ is sampled from an exponential distribution in the case of kinetic Monte Carlo dynamics.
The partial evolution scheme is related to patch dynamics in the equation-free framework (EFF) for multiscale simulation [24, 28, 41], in that both approaches evolve the microscopic dynamics on small spatial patches for short time intervals. Despite this similarity, we adopt it for a different purpose. The EFF makes use of patch dynamics as a simulation tool to advance macroscopic observables through repeated microscopic simulations, without explicitly deriving macroscopic dynamical equations. In contrast, we use partial evolution scheme as a data-generation mechanism for learning macroscopic dynamics.
We note that for the partial evolution scheme to be accurate, the microscopic dynamics within a small patch must evolve as if embedded in the full system. To achieve this, when evolving the microscopic dynamics within a local patch, appropriate boundary conditions should be imposed. In our implementation, the local environment is handled in two possible ways. One option is to treat the neighboring sites outside the patch as ghost cells, whose values are held fixed during the short-time evolution. This approach does not increase the computational cost of the partial evolution scheme. Alternatively, one may introduce a thin buffer region surrounding the patch. In this approach, the patch and buffer evolve simultaneously, but only the central states are retained for the macroscopic derivation. The purpose of the buffer is to shield the interior dynamics from artificial boundary effects. This strategy is commonly used in the patch dynamics [45, 47, 46]. While this strategy slightly increases the computational cost, we will keep the buffer thin relative to the full system size, ensuring the cost of the partial evolution scheme remains low.
We will demonstrate how to derive the macroscopic dynamics from the training data pairs obtained from the partial evolution scheme.
@@
Once the autoencoder is trained, we will generate the latent training data pairs $\{\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}}\}$ for the macroscopic dynamics derivation:
@@
where $\mathbf{x}_{t,\mathcal{I}}$ denotes the restriction of $\mathbf{x}_{t}$ to the local patch $\mathcal{I}$. Next, we introduce the process of deriving macroscopic dynamics.
Macroscopic dynamics derivation.
@@
where $\boldsymbol{\mu}$ is the drift term and $\boldsymbol{\Sigma}$ is the diffusion term. In most experiments, we adopt fully connected networks for both $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$. For the NbMoTa alloy experiment in Section III.3, we adopt OnsagerNet [57, 12] for $\boldsymbol{\mu}$ to account for more complex macroscopic dynamics. The OnsagerNet can capture physically interpretable and stable macroscopic dynamics by incorporating the generalized Onsager principle into the model structure, and its form is given by:
$\boldsymbol{\mu}(\mathbf{z}_{t})=-(\mathbf{M}(\mathbf{z}_{t})+\mathbf{W}(\mathbf{z}_{t}))\nabla V(\mathbf{z}_{t})+\mathbf{f}(\mathbf{z}_{t}),$ (5)
@@
Existing works train the SDE by minimizing the negative log-likelihood [12, 19, 60, 26]:
$\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{z}_{t},\mathbf{z}_{t+\delta t}}[-2\log$ (6)
$\quad p\left(\mathbf{z}_{t+\delta t}\,\big|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t\right)],$
@@
$\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}}}[-2\log$ (7)
$\quad p(\mathbf{z}_{t+\delta t,\mathcal{I}}\,|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,K\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t)].$
The only difference between $\mathcal{L}$ and $\mathcal{L}_{p}$ is that the covariance term in $\mathcal{L}_{p}$ is multiplied by a factor $K$, where $K$ denotes the number of patches introduced earlier. We can interpret the influence of $K$ qualitatively. The factor $K$ in the loss compensates for the additional stochasticity, leading to a smaller learned diffusion term. During the data generation of $\mathbf{x}_{t+\delta t,\mathcal{I}}$ from $\mathbf{x}_{t}$, we introduce additional randomness by performing partial evolution. Therefore, in the derivation of macroscopic dynamics, we multiply the diffusion term by $K$ to correct for the extra stochasticity.
@@
It is important to observe that in Eq. 3, we define $\mathbf{z}_{t+\delta t,\mathcal{I}}$ in a nontrivial way, while the most naive way will be $\mathbf{z}_{t+\delta t,\mathcal{I}}=\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})$. Note that the loss function $\mathcal{L}_{p}(\boldsymbol{\theta})$ involves the weighted norm of the residual $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}-\boldsymbol{\mu}(\mathbf{z}_{t})\delta t$ since $p$ represents a Gaussian distribution. If we define $\mathbf{z}_{t+\delta t,\mathcal{I}}=\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})$ in the naive way, we will have:
@@
which is the average of $\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})-\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{J}})$. Since $\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})$ may be very different from $\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{J}})$ when $\mathcal{J}\neq\mathcal{I}$, the resulting $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}$ will be very noisy. However, the definition in Eq. 3 will yield:
@@
which directly measures the change in the patch $\mathcal{I}$, thus reducing noise and leading to more stable training. We refer to the approach that adopts the naive formulation $\mathbf{z}_{t+\delta t,\mathcal{I}}=\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})$ and derives the macroscopic dynamics via $\mathcal{L}$ as the baseline. A detailed comparison of our method with the baseline is provided in Section III.
@@
In many microscopic systems, the microscopic interactions between lattice sites are local. For instance, microscopic interactions are limited to the first nearest neighbor in the Ising model, and are limited to a finite cutoff distance in alloy systems. Under this locality assumption, when the time increment $\delta t$ is sufficiently small, the state increments on disjoint spatial patches are approximately independent. Specifically, define the short-time increments
$\Delta\mathbf{z}_{t}\coloneqq\mathbf{z}_{t+\delta t}-\mathbf{z}_{t},$ $\Delta\mathbf{z}_{t}^{\ast}\coloneqq\mathbf{z}^{\ast}_{t+\delta t}-\mathbf{z}^{\ast}_{t},$ (11)
$\Delta\mathbf{z}_{t,\mathcal{I}}\coloneqq\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t},$ $\Delta\mathbf{z}^{\ast}_{t,\mathcal{I}}\coloneqq\mathbf{z}^{\ast}_{t+\delta t,\mathcal{I}}-\mathbf{z}^{\ast}_{t}.$
Then, for two disjoint patches $\mathcal{I}\neq\mathcal{J}$, we have
$\Delta\mathbf{z}_{t,\mathcal{I}}$ $=\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{I}})-\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{I}}),$ (12)
$\Delta\mathbf{z}_{t,\mathcal{J}}$ $=\boldsymbol{\varphi}(\mathbf{x}_{t+\delta t,\mathcal{J}})-\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{J}}),$
which can be treated as approximately independent random variables.
Let $\hat{q}(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})$ denote the distribution of $\mathbf{x}_{t+\delta t,\mathcal{I}}$ derived by evolving the full system from $\mathbf{x}_{t}$ for a time step $\delta t$ and subsequently restricting the state to the local patch $\mathcal{I}$. Under the local interaction assumption and for sufficiently small $\delta t$, the distribution $q$ can be well approximated by the partial evolution distribution $\hat{q}$:
$q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})\approx\hat{q}(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}).$ (13)
Furthermore, since we consider a sufficiently large microscopic system and the macroscopic observables are intensive quantities, the macroscopic observable of the full system can be approximated by the average over local patches: